OpenStudy (anonymous):

PleAe help the area of a window is 18 ft squared the length of the window is two times the width of the window what are the demensions of the window

OpenStudy (anonymous):

\[x\times 2x=18\] \[2x^2=18\] \[x^2=9\] \[x=3\] so dimenstions are 3 and 6

OpenStudy (anonymous):

l = 2w area = l * w = 2w * w = 2w^2 = 18 => w^2 = 9 => w = 3

OpenStudy (anonymous):

l = 2w = 6